Mary is twice as old as John and half as old as Bob.in 22 years time, Bob will be twice as old as John. how old is Mary?
Found 2 solutions by rwm, mananth:
Answer by rwm(914)

m=2j
m=b/2
(b+22)=2*(j+22)
b=44
j=11
m=22

Answer by mananth(16949)

Mary =x
John= x/2 (because mary is twice Bob's age)
Bob = 2x ( mary is 1/2 John's age)
After 22 years
Bob = 2x+22
John = x/2 + 22
2x+22= 2(x/2+22)
2x+22=x+44
2x-x=44-22
x=22 Mary's age